Main INDEX
Monthly INDEX
PREV
NEXT
User name Emily
Log entry time 01:20:09 on August21,2006
Entry number 181231
keyword=Swing Shift/Run Summary
Swing Shift Summary
16:00 - On Kin point 38, finishing up Rate Check with LD2 and radiator out.
16:41 - VESDA alarm: In zone, Urgent, Air Flow. In Beamway. Called MCC.
They saw the problem too and came over and reset it.
16:50 - Asked for 11uA. Put radiator in 5%. Started last part of rate check
17:25 - MCC took beam for 15min to do tests.
17:40 - Beam is back...Started Gamma-D production runs
LD2, 5% radiator, 11uA.
18:32 - Q1 died. Tried to set to 0amps, turn off, reset, turn on. Did nothing.
Pressed the green button #9 in the middle room, and tried resetting
again. Now it works. Took ~10min to fix, all together.
19:48 - Q1 died again. Reset and used green button. Fixed in 8min.
20:47 - Q1 had yet another Earth Leakage Failure. Green button 4 times.
Fixed in 10min.
MCC called saying we are authorized to run with one ion chamber masked
because the left and right chambers are redundant.
22:50 -
Many Magnet Problems:
LQ1: Same problems as before. Earth Leakage Failure, Max Current Set, Spare Interlock, and it stopped reading back. Fixed it using green button pushing method, and on gui. Took a few tries for it to work.
LQ2: This one is fine.
LD1: This does have power supply on, and it looks fine in the LD1 System GUI (reads back a constant 50.1) but on the Tools GUI there is no readback, Had a Lead A Flow fault and a Dump Mode Fast that was red, but these dissapeared when the reset button in the GUI was pressed.
LQ2: This one is fine.
LD1: This does have power supply on, and it looks fine in the LD1 System GUI (reads back a constant 50.1) but on the Tools GUI there is no readback, Had a Lead A Flow fault and a Dump Mode Fast that was red, but these dissapeared when the reset button in the GUI was pressed.
LQ3: There was an Interlock fault which went away when reset was pressed, but the Quench fault remained. It won't go away just resetting in the GUI.
RQ1 is fine.
RQ2: There was an Interlock fault which went away when reset was pressed, but the Quench fault remained. It won't go away just resetting in the GUI.
RD1: There were faults in the Lead A Flow and Lead B Flow but they went away when reset was pressed in the GUI. There is a Power Supply fault that goes away when you press the "on" button, but then it turns itself off and the fault returns (about 2 seconds later). It is Off and Disengaged.
RQ3: There was an Interlock fault which went away when reset was pressed, but the Quench fault remained. It won't go away just resetting in the GUI.
ALSO: The N2 level CLL71Q210 is around 23 or 24, which is under the required 25. The rest of these and the H2 levels look fine.
Went down to the hall with Rob Feuerbach. Reset magnets and rebooted IOCs. Eventually things looked better.
1:00 - On the Left HRS, all magnets are now responding as they should. Didn't touch the RHRS magnets...they're not as important right now.
Swing Run Summary
Run Num, Num Events, Tot Current
2993/21302 869K, 1.2m, 20uA
2998/21304 853K, 313k
2999/21305 1.9m, 1.0m
3000/21306 1.9m, 1.0m
3001/21307 1.9m, 1.0m
3002 JUNK
3003/21309 1.9m, 1.0m
3004/21310 1.9m, 1.0m
3005/21311 1.8m, 997m
3006/21312 1.9m, 1.0m
3007/21313 1.9m, 1.0m
3008/21314 1.9m, 1.0m
3009/21315 1.9m, 1.0m
3010/21316 1.9m, 1.0m
3011/21317 1.9m, 1.0m
3012/21318 1.9m, 1.0m
3013/21319 1.9m, 1.0m
3014/21320 1.9m, 1.0m
3015/21321 1.2m, 0.65m